Given the function F(w,x,y,z)=Σ(1,3,7,11,15)+dc(0,2,5,8)
i. Write the function in conjunctive normal form [3 Marks]
ii. Minimize the function (DNF) using Karnaugh Map [7 Marks]
1
Expert's answer
2020-06-25T08:08:02-0400

(i)

In conjuctive normal form

The function is represented in POS form as product of sum

There for the given function

F(w,x,y,z)= (1,3,7,11,15)+dc(0,2,5,8)

Fcnf= (z'+w'+x').(w+y'+z).(z'+w)

(ii)

In disjunctive normal form

The function is represented in SOP form as sum of product

The given function is

F(w,x,y,z)= (1,3,7,11,15)+dc(0,2,5,8)

Fdnf = z(y+(y+w)')